A steel rod. 31 cm. long is to be cut into 3 parts, two of which parts are to be of equal lengths. The lengths of each of the 2 equal parts is 5 cm. more than the third part. Find their lengths.
Found 2 solutions by longjonsilver, askmemath:
Answer by longjonsilver(2297)

again, just build it up gradually. Start with a definition.

Let x = length of "third" piece

so, each longer piece is x+5

so we have x+5 + x+5 + x = 31
3x + 10 = 31
3x = 21
x = 21/3
x = 7

So, pieces are 12, 12 and 7cm long

jon.

Answer by askmemath(368)

Let the third part be of length l
Then the length of the 2 unequal parts which are greater than the third part by 5 cms is l+5
Total lenght = 31
l+5+l+5+l=31
3l+10=31
Subtracting 10 on both sides
3l=31-10
3l=21
Dividing by 3 throughout
l=7
And so the other parts are of length 12cms each.